Y=X+4X2-0.2X3 (Y=is wheat fields in kgs and X=amount of ferterlizer applied) Derive MPP and APP of the variable factor


1
Expert's answer
2021-05-13T10:03:31-0400

MPP "= \frac{dY}{dx} = 1 + 8x - 0.6x^{2}"

APP "= \frac{Y}{x} = 1 + 4x -0.2x^{2}"


MPP is maximum when "\frac{\delta MPP}{\delta x} = 0"

"8-1.2x = 0"

"1.2x = 8"

"x = 6.67"


APP is maximum when "\frac{\delta APP}{\delta x} = 0"

"4-0.4x = 0"

"0.4x = 4"

"x = 10"